OpenStudy (anonymous):

a regular polygon with an interior angle sum of 9000°? If so, what is it?

OpenStudy (anonymous):

9000= (n-2)180 (n number of sides)

OpenStudy (anonymous):

N= number of sides right?

OpenStudy (anonymous):

using the formula sum of all the interior angles= (n-2)*180 degree where n = number of sides of the polygon hence 9000 = (n-2)180 divide both sides by 180 \[9000/180 = n-2\] \[50 = n-2\] transpose -2 to right hand side \[50+2=n\] \[52=n\] hence it is a 50-gon

OpenStudy (anonymous):

Um.....

OpenStudy (anonymous):

what does transpose mean?

OpenStudy (anonymous):

of a matrix?

OpenStudy (anonymous):

I dont know, it the answer the kushashwa gave, he says to "transpose the -2 to the right hand side"

OpenStudy (anonymous):

It's just short for add 2 to both sides....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I was looking over your method

OpenStudy (anonymous):

9000/180 = number of lines???????

OpenStudy (anonymous):

? 9000 = (n-2)180 = 180n-360 -> 180 n = 9360 -> n= 9360/180 = 52

OpenStudy (anonymous):

maybe I don't understand the symbols and I am sorry if this is a silly question, but now did you go from 180n to 9360??

OpenStudy (anonymous):

"transpose" the 360 from right side to left side (ie add 360 to both sides).